OpenStudy (anonymous):

You deposit $500 into a savings account that earns 3.2% annual interest. How much is in the savings account after 10 years?

OpenStudy (jdoe0001):

\(\bf \textit{simple interest}=deposit\cdot rate\textit{(in decimal format)}\cdot \textit{time(years)}\\ \quad \\\implies i = prt\\ \quad \\ rate = 3.2\%\implies 0.032\qquad thus\\ \quad \\ i=500\cdot 0.032\cdot 10\)

OpenStudy (jdoe0001):

so the savings will have the original deposit PLUS the earned "interest"